A pet store reported the following: Net Sales.. $689,000 Cost of Goods Sold.. $357,000 Operating Expenses.. $202,000 What is the “$ Net Profit/Net Loss” and “% for Operating Expenses” (round to one decimal place)?

A. $130,000 Net Profit; 29.3% Operating Expenses

B. $130,000 Net Loss; 29.3% Operating Expenses

C. $534,000 Net Loss; 77.5% Operating Expenses

D. $534,000 Net Profit; 77.5% Operating Expenses
Answer #1

Net profit=Net Sales-COGS-Operating Expenses

=(689000-357000-202000)=$130000

Operating expenses%=Operating expenses/Net Sales

=(202000/689000)=29.3%(Approx).

Hence the correct option is A.